In order to help you through college, your parents just deposited $20,000 into a bank account paying 6% interest.

Starting tomorrow, you plan to withdraw equal amounts from the account at the beginning of each month for the next four years.

What is the most you can withdraw annually?
Answer #1

Ans $ 5445.12

Annuity PV Due (Beginning)= P + P [ 1 - ( 1 + r )-(n-1) ]
                       r
20000 = P + P * ( 1 - ((1 / (1 + 6%)^(4-1))))/ (6%)
20000 = P + 2.67301194946164 P
20000 = 3.67301194946164 P
P = 20000 / 3.67301194946164
P = 5445.12